OpenStudy (anonymous):

Solve this inequality: |x – 4 | ≤ 6.

OpenStudy (anonymous):

Absolute would tune x-4 to x+4. Solve for x.

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

<<

OpenStudy (anonymous):

\[|x-4|\leq 6\\ -6\leq x-4\leq 6\] add \(4\) all the way across

OpenStudy (anonymous):

I can't beleive I forgot that, omg , im stupid

OpenStudy (bradely):

|x – 4 | ≤ 6. x-4≤ +6 or -6 source : http://www.mathskey.com/question2answer/ask